Haunted Forest, Inc.is selling fog machines.

Use the following information about Haunted Forest, Inc. to answer the following questions.

Average selling price per unit $328.

Variable cost per unit $208

Units sold 345

Fixed costs $15,907

Interest expense $3,001

Based on the data above, what is the degree of total (combined) leverage of Haunted Forest, Inc.?

Round the answer to two decimals
Answer #1
1.84
Dear Student
Thank you for referring Chegg
Please give thumbs up if you are satisfied
Statement showing Computations
Particulars Amount
Sales = 345*328            113,160.00
Less Variable Expenses= 345 * 208            (71,760.00)
Contribution Margin = 113,160 - 71,760               41,400.00
Less Fixed cost            (15,907.00)
EBIT               25,493.00
Less Interest expense               (3,001.00)
EBT               22,492.00
degree of total (combined) leverage = Contribution / EBT = 41,400/22,492                         1.84